Calculate the rate constant,
k
, for a reaction at
66.0
°C that has an activation energy of
78.5
kJ/mol and a frequency factor of
5.67
×
10
11
s−1.
k
=

s

1
1
Expert's answer
2020-04-06T11:44:06-0400
"k = Ae^{{-E_a \over RT}}"

"k = 5.67*10^{-11}*2.71^{-78.5 \over 8.314*339.15} = 5.83*10^{-11}s^{-1}"